OpenStudy (anonymous):

Use the steps listed in the lesson to solve the following: 3( x - 4 ) + 2 = 23 Each step must be shown and explained.

OpenStudy (compassionate):

3x - 12 + 2 = 23 - Distribute 3x - 10 = 23 - add like terms 3x + 10 - 10 = 23 + 10 - Do the opposite and add ten 3x = 33 -- --- - Divide both sides by 3 3 3 x = ? What is 33/3? You do the last step

OpenStudy (anonymous):

33/3=11